OpenStudy (sf_r):

What are the zeros of the polynomial equation f(x)=x^2+4x-60

rishavraj (rishavraj):

factorise it....

OpenStudy (dayakar):

LET f(x) =0 x^2+4x-60 =0 x^2 +10x -6x -60=0 [ splitting the middle term] can u find remaining part

OpenStudy (sf_r):

Why did you split the middle term?

OpenStudy (dayakar):

to get common factors

rishavraj (rishavraj):

@SF_R http://www.purplemath.com/modules/factquad.htm

OpenStudy (sf_r):

They are no common factors

OpenStudy (sf_r):

Do I use the Quadratic method?

OpenStudy (dayakar):

x is common in first two terms and -6 is common in next two terms

OpenStudy (sf_r):

I am pretty sure they have to be common in all terms

OpenStudy (dayakar):

yes ,of course u can shall i complete this before that

OpenStudy (sf_r):

I have the answer choices A. -10,6 B.-14, 5 C. -6, 10 D. -5,14

OpenStudy (dayakar):

\[x ^{2}+10x-6x-60=0\] x*x +10x - 6x - 6*10 =0 x(x+10) -6(x+10) =0 (x+10) (x-6) =0 x+10 =0 or x-6 =0 now can u find x

OpenStudy (dayakar):

u go through the solution

OpenStudy (sf_r):

now you minus 10 and 0 and same for the other side but we are going to add

OpenStudy (dayakar):

yes, write the answer

OpenStudy (sf_r):

-10, 6

OpenStudy (dayakar):

exactly,

OpenStudy (dayakar):

do u understand

OpenStudy (sf_r):

Yes! Thank you so much! :)